### v1

# Sequential FDR and pFDR control under arbitrary dependence, with application to pharmacovigilance database monitoring

Abstract
We propose sequential multiple testing procedures which control the false discover rate (FDR) or the positive false discovery rate (pFDR) under arbitrary dependence between the data streams. This is accomplished by “optimizing” an upper bound on these error metrics for a class of step down sequential testing procedures. Both open-ended and truncated versions of these sequential procedures are given, both being able to control both the type I multiple testing metric (FDR or pFDR) at specified levels, and the former being able to control both the type I and type II (e.g., FDR and the false nondiscovery rate, FNR). In simulation studies, these procedures provide 45-65% savings in average sample size over their fixed-sample competitors. We illustrate our procedures on drug data from the United Kingdom’s Yellow Card Pharmacovigilance Database.
1 Introduction
The majority of the procedures proposed in the statistics literature for multiple testing for fixed sample size or sequential data can be bifurcated into either step up or step down procedures. Step up procedures decide whether to accept or reject null hypotheses in order of increasing significance, whereas step down procedures operate in the reverse order. For testing $J\geq 2$ null hypotheses, the Benjamini-Hochberg (1995, hereafter BH) procedure is step up, and is known to control the false discovery rate (FDR) at the nominal level $\alpha$ under independence of $p$-values, and at no more than the inflated level $C\alpha$, where $C=C(J):=\sum_{j=1}^{J}1/j$, under arbitrary dependence among $p$-values, with $C\approx\log(J)$ for large $J$. Although Benjamini and Yekutieli, (2001) showed that FDR can be controlled at the uninflated level $\alpha$ under certain types of positive dependence, Guo and Rao, (2008) showed that the FDR bound $C\alpha$ is sharp for the step up BH procedure in the sense that there exists a joint distribution of $p$-values for which the FDR equals this. Analogous results were established by Bartroff, (2018) and Bartroff and Song, (2020) about step up procedures for sequential data.
In this sense, the “worst case” performance of step up procedures is essentially completely understood, at least without imposing additional restrictions or assumptions on the joint distribution of $p$-values. So in this paper we turn out attention to studying and controlling the worst case performance of step down procedures, in the setting of sequential data, i.e., $J$ data streams or arbitrary dependence. More specifically, we give a step down procedure for sequential data and the smallest possible constant $D<C$ whose FDR is bounded above by $D\alpha$ where, again, $\alpha$ is the nominal FDR level.
Hart and Weiss, (1997) used linear programming to find a sharp upper bound on the probability of rejection for a class of multiple testing procedures in terms of the true, marginal c.d.f.s of the $p$-values. Calculating their bound therefore requires knowledge of at leas the marginals of the true joint distribution. Guo and Rao, (2008) also used optimization but to find an upper bound on the false discovery rate (FDR) in terms of a step down procedure’s step values, that holds under arbitrary joint distribution of the $p$-values. By choosing the step values to control the upper bound, they produce a step down procedure whose FDR control holds under arbitrary joint distributions. The current paper extends this line of research in the following ways. First, by extending to the sequential setting an FDR controlling procedure without requiring or imposing assumptions on the joint distribution of the $p$-values; the open-ended versions of these sequential procedures allow the possibility of simultaneously controlling both type I and type II error metrics (e.g., FDR and the false nondiscovery rate, FNR). Secondly, by removing an implicit assumption on the $p$-values in Guo and Rao’s proof that is stronger than validity of $p$-values and applies to both the sequential and fixed-sample settings; see Appendix A for details. Finally, by extending this methodology to positive false discovery rate (pFDR) and its type II analog.
The rest of this paper is organized as follows. After giving the notational set up in Section 2, Section 3 covers untruncated (i.e., open-ended) sequential procedures and records the FDR/FNR- and pFDR/pFNR-controlling procedures in Theorems 3.1 and 3.2, respectively, as special cases of a generic untruncated procedure given in Section 3.1. Truncated sequential procedures are handled analogously in Section 4 with the corresponding FDR- and pFDR-controlling procedures given in Theorems 4.1 and 4.2, respectively. To aide the reader, Table 1 is a “directory” of these theorems and cases. In Section 5 we present simulation studies on Binomial and Poisson data of the proposed sequential procedures and compare with their fixed sample counterparts, and in Section 6 we apply our proposed procedures to data from the UK’s Yellow Card Pharmacovigilance Database. We end with conclusions and discussion in Section 7. Proofs are given in the Appendix.
Table 1: Directory of this paper’s error control theorems
Error Metrics Untruncated, Type I and II Controlling Truncated, Rejective
[…]
Dedication
This paper is dedicated to the memory of Prof. Tze Leung Lai, who influenced and inspired the authors directly and indirectly. While the current paper is not a direct offshoot of one of Prof. Lai’s works, it overlaps with some of Lai’s most active research areas including sequential hypothesis testing (Lai,, 1981, 1988, 2001), adaptive designs (Lai and Robbins,, 1979; Lai,, 1987), multiple comparisons (Lai,, 2000; Bartroff and Lai,, 2010), optimization (Han et al.,, 2006), quality control (Lai,, 1995), longitudinal data (Lai et al.,, 2006), and biomedical applications (Bartroff et al.,, 2013), to name a few references.
2 Set up
2.1 Data Streams, Hypotheses, and Error Metrics
Our general form of data streams and hypotheses is as follows. Assume that there are $J\geq 2$ data streams
Data stream $1$: $X_{1}^{(1)},X_{2}^{(1)},\ldots$
Data stream $2$: $X_{1}^{(2)},X_{2}^{(2)},\ldots$ (1)
$\vdots$
Data stream $J$: $X_{1}^{(J)},X_{2}^{(J)},\ldots.$
In order to implement our proposed procedures, the marginal type I and II error probabilities of component test statistics on each stream will need to be controlled controlled at certain levels; see (4)-(5) below. Beyond that, we make no assumptions about the dimension of the sequentially-observed data $X_{n}^{(j)}$, which may themselves be vectors of varying size, nor about the dependence structure of within-stream data $X_{n}^{(j)},X_{m}^{(j)}$ or between-stream data $X_{n}^{(j)},X_{m}^{(j^{\prime})}$ ($j\neq j^{\prime}$). In particular there can be arbitrary “overlap” between data streams, an extreme case being that all the data streams are the same, which is equivalent to testing multiple hypotheses about a single data source. For any positive integer $j$ let $[j]=\{1,\ldots,j\}$. For each data stream $j\in[J]$, assume that there is a parameter vector $\theta^{(j)}\in\Theta^{(j)}$ determining that distribution of the stream $X_{1}^{(j)},X_{2}^{(j)},\ldots$, and it is desired to test a null hypothesis $H^{(j)}$ versus the alternative hypothesis $G^{(j)}$, where $H^{(j)}$ and $G^{(j)}$ are disjoint subsets of the parameter space $\Theta^{(j)}$ containing $\theta^{(j)}$. The null $H^{(j)}$ is considered true if $\theta^{(j)}\in H^{(j)}$, and false if $\theta^{(j)}\in G^{(j)}$. The global parameter $\theta=(\theta^{(1)},\ldots,\theta^{(J)})$ is the concatenation of the individual parameters and is contained in the global parameter space $\Theta=\Theta^{(1)}\times\cdots\times\Theta^{(J)}$.
The general notation (1) includes fully-sequential sampling where the streamwise sample sizes may take any value $1,2,\ldots$ ad infinitum, but other set ups including group sequential, truncated, and even fixed sample size can be included in this as well. For example, the $n$th “observation” $X_{n}^{(j)}$ in the $j$th stream may actually be the $n$th group $X_{n}^{(j)}=(X_{n,1}^{(j)},\ldots,X_{n,\ell}^{(j)})$ of size $\ell$. Moreover, the group size $\ell$ may vary with $n$ and may even be data-dependent, e.g., determined by some type of adaptive sampling. Similarly, truncated sequential (or group sequential) sampling can be implemented for the $j$th stream by defining $X_{n}^{(j)}=\emptyset$ for all $n>\overline{N}^{(j)}$ for some stream-specific truncation point $\overline{N}^{(j)}$, or globally for all streams by replacing statements like “for some $n$” in what follows with “for some $n\leq\overline{N}$,” for some global truncation point $\overline{N}$. One may represent a fixed sample size in the $j$ the stream in this way by taking $\overline{N}^{(j)}=1$, or in all streams with $\overline{N}=1$.
For any multiple testing procedure under consideration, let $V$ denote the number of true null hypotheses it rejects (i.e., the number of false positives), $W$ the number of false null hypotheses it accepts (i.e., the number of false negatives), and $R$ the number of null hypotheses it rejects. The number of null hypotheses accepted is therefore $J-R$. Under true value of the parameter $\theta$, the false discovery and non-discovery rates (Benjamini and Hochberg,, 1995, FDR, FNR) are
$\mbox{FDR}=\mbox{FDR}(\theta)=E_{\theta}\left(\frac{V}{R\vee 1}\right),\quad\mbox{FNR}=\mbox{FNR}(\theta)=E_{\theta}\left(\frac{W}{(J-R)\vee 1}\right),$
where $x\vee y=\max\{x,y\}$. Similarly, the positive false discovery rate pFDR (Storey,, 2002) and its type II analog the positive false nondiscovery rate pFNR are defined as
$\mbox{pFDR}=\mbox{pFDR}(\theta)=E_{\theta}\left(\left.\frac{V}{R}\right|R\geq 1\right)\quad\mbox{and}\quad\mbox{pFNR}=\mbox{pFNR}(\theta)=E_{\theta}\left(\left.\frac{W}{J-R}\right|J-R\geq 1\right).$ (2)
2.2 Test Statistics and Critical Values
The building blocks of the sequential procedures defined below are $J$ individual sequential test statistics $\{\Lambda^{(j)}(n)\}_{j\in[J],\;n\geq 1}$, where $\Lambda^{(j)}(n)$ is the statistic for testing $H^{(j)}$ vs. $G^{(j)}$ based on the data $X_{1}^{(j)},X_{2}^{(j)},\ldots,X_{n}^{(j)}$ available from the $j$th stream at time $n$. For example, in parametric settings $\Lambda^{(j)}(n)$ may be a sequential log (generalized) likelihood ratio statistic for testing $H^{(j)}$ vs. $G^{(j)}$. Like the fixed sample size procedures mentioned above, our sequential procedures will utilize step values (or step value vectors) which are $J$-long vectors of non-decreasing values in $[0,1]$, such as
$\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{J})\quad\mbox{with}\quad 0<\alpha_{1}\leq\alpha_{2}\leq\ldots\leq\alpha_{J}\leq 1.$ (3)
Given two step value vectors $\bm{\alpha}$ and $\bm{\beta}$, we say that a sequential test statistic $\{\Lambda^{(j)}(n)\}_{n\geq 1}$ is implemented with step values $\bm{\alpha},\bm{\beta}$ if there are critical values $\{A_{k}^{(j)},B_{k}^{(j)}\}_{k\in[J]}$ such that
$P_{\theta^{(j)}}(\Lambda^{(j)}(n)\geq B_{k}^{(j)}\;\mbox{some $n$,}\;\Lambda^{(j)}(n^{\prime})>A_{1}^{(j)}\;\mbox{all $n^{\prime}<n$})$ $\leq\alpha_{k}\quad\mbox{for all}\quad\theta^{(j)}\in H^{(j)}$ (4)
$P_{\theta^{(j)}}(\Lambda^{(j)}(n)\leq A_{k}^{(j)}\;\mbox{some $n$,}\;\Lambda^{(j)}(n^{\prime})<B_{1}^{(j)}\;\mbox{all $n^{\prime}<n$})$ $\leq\beta_{k}\quad\mbox{for all}\quad\theta^{(j)}\in G^{(j)}$ (5)
for all $k\in[J]$. The critical values $A_{1}^{(j)},B_{1}^{(j)}$ are simply the critical values for the sequential test that samples until $\Lambda^{(j)}(n)\not\in(A_{1}^{(j)},B_{1}^{(j)})$, and this test has type I and II error probabilities bounded above by $\alpha_{1}$ and $\beta_{1}$, respectively. The values $B_{k}^{(j)}$, $k\in[J]$, are then such that the similar sequential test with critical values $A_{1}^{(j)}$ and $B_{k}^{(j)}$ has type I error probability $\alpha_{k}$, which is just a restatement of (4), with an analogous statement holding for critical values $A_{k}^{(j)}$ and $B_{1}^{(j)}$, type II error probability $\beta_{k}$, and (5). We say that the test statistics $\{\Lambda^{(j)}(n)\}_{j\in[J],\;n\geq 1}$ for all the streams are implemented with step values $\bm{\alpha},\bm{\beta}$ if they are for each stream. In all commonly-encountered testing situations there are standard sequential statistics which can be implemented with given step values $\bm{\alpha},\bm{\beta}$. Bartroff and Song, (2014, 2020) and Bartroff, (2018) give examples.
Without loss of generality we assume that, for each $j\in[J]$,
$A_{1}^{(j)}\leq A_{2}^{(j)}\leq\ldots\leq A_{J}^{(j)}\leq B_{J}^{(j)}\leq B_{J-1}^{(j)}\leq\ldots\leq B_{1}^{(j)},$ (6)
$A_{k}^{(j)}=A_{k+1}^{(j)}$ if and only if $\beta_{k}=\beta_{k+1}$ (7)
$B_{k}^{(j)}=B_{k+1}^{(j)}$ if and only if $\alpha_{k}=\alpha_{k+1}$. (8)
The sequential multiple testing procedures proposed below will involve ranking the test statistics associated with different data streams, which may be on completely different scales in general, so for each stream $j$ we introduce a standardizing function $\varphi^{(j)}(\cdot)$ which will be applied to the statistic $\Lambda^{(j)}(n)$ before ranking. The standardizing functions $\varphi^{(j)}$ can be any increasing functions such that $\varphi^{(j)}(A_{k}^{(j)})$ and $\varphi^{(j)}(B_{k}^{(j)})$ do not depend on $j$, and we let
$a_{k}=\varphi^{(j)}(A_{k}^{(j)})\quad\mbox{and}\quad b_{k}=\varphi^{(j)}(B_{k}^{(j)}),\quad j,k\in[J],$ (9)
denote these common values. Given critical values $\{A_{k}^{(j)},B_{k}^{(j)}\}_{j,k\in[J]}$ satisfying (4)-(5), one may choose arbitrary values $\{a_{k},b_{k}\}_{k\in[J]}$ satisfying the same monotonicity conditions as the $\{A_{k}^{(j)},B_{k}^{(j)}\}$ according to (7)-(8) and then define the standardizing functions $\varphi^{(j)}(\cdot)$ to be increasing, piecewise linear functions satisfying (9). For example, if all the $\alpha_{k}$ are distinct and the $\beta_{k}$ are distinct then a simple choice for the $\{a_{j},b_{j}\}$ are the integers
$a_{1}=-J,\quad a_{2}=-J+1,\quad\ldots,\quad a_{J}=-1,\quad b_{J}=1,\quad b_{J-1}=2,\quad\ldots,\quad b_{1}=J.$
In any case, the assumptions on the critical values and standardizing functions imply that the $a_{k}$ must be nondecreasing and the $b_{k}$ nonincreasing. Finally, we denote $\widetilde{\Lambda}^{(j)}(n)=\varphi^{(j)}(\Lambda^{(j)}(n))$ and then (4)-(5) can be written
$P_{\theta^{(j)}}(\widetilde{\Lambda}^{(j)}(n)\geq b_{k}\;\mbox{some $n$,}\;\widetilde{\Lambda}^{(j)}(n^{\prime})>a_{1}\;\mbox{all $n^{\prime}<n$})$ $\leq\alpha_{k}\quad\mbox{for all}\quad\theta^{(j)}\in H^{(j)}$ (10)
$P_{\theta^{(j)}}(\widetilde{\Lambda}^{(j)}(n)\leq a_{k}\;\mbox{some $n$,}\;\widetilde{\Lambda}^{(j)}(n^{\prime})<b_{1}\;\mbox{all $n^{\prime}<n$})$ $\leq\beta_{k}\quad\mbox{for all}\quad\theta^{(j)}\in G^{(j)},$ (11)
for all $j,k\in[J]$.
The theorems in this paper will give upper bounds on the FDR and pFDR, and their type II, of versions sequential procedures implemented with arbitrary step values $\bm{\alpha},\bm{\beta}$. In practice we recommend using some of the commonly used values such as those recommended by Benjamini and Hochberg, (1995) or Benjamini and Liu, (1999). For desired value $q_{1}\in(0,1)$ of the type I error metric (FDR or pFDR), the former are
$\bm{\alpha}_{j}=\bm{\alpha}(q_{1})_{j}:=q_{1}j/J\quad\mbox{for}\quad j\in[J],$ (12)
and the latter are
$\bm{\alpha}_{j}=\bm{\alpha}(q_{1})_{j}=1-\left(1-\left(1\wedge\frac{q_{1}J}{J-j+1}\right)\right)^{1/(J-j+1)}\quad\mbox{for}\quad j\in[J].$ (13)
Here $x\wedge y=\min\{x,y\}$. We recommend the using the same expressions for $\bm{\beta}=\bm{\beta}(q_{2})$ for desired value $q_{2}\in(0,1)$ of the type II error metric (FNR or pFNR). In our examples below we will refer to (12) and (13) as the BH and BL step values, respectively.
2.3 Simple hypotheses and Wald approximations
In practice, many testing situations can be reduced to, or approximated by, testing simple vs. simple hypotheses. Müller et al., (2007, Section 1) point out that testing a battery of simple-vs.-simple hypothesis tests has been the standard set up for use of FDR in the literature. In this section we show how to construct the test statistics $\Lambda^{(j)}(n)$ and critical values $A_{s}^{(j)},B_{s}^{(j)}$ satisfying (4)-(5) for any data stream $j$ such that $H^{(j)}$ and $G^{(j)}$ are both simple hypotheses. In this case the test statistic $\Lambda^{(j)}(n)$ can be taken to be log-likelihood ratio (density under $G^{(j)}$ divided by the density under $H^{(j)}$) yielding the sequential probability ratio test (SPRT; see Chernoff,, 1972). See Bartroff and Song, (2020, Section 5.1) for a more formal description of hypotheses and parameter space in this set up. In the single null hypothesis testing set up, the SPRT samples until $\Lambda^{(j)}(n)\not\in(A,B)$ where the critical values $A,B$ are chosen so that
$P_{H^{(j)}}(\Lambda^{(j)}(n)\geq B\;\mbox{some $n$,}\;\Lambda^{(j)}(n^{\prime})>A\;\mbox{all $n^{\prime}<n$})$ $\leq\alpha$ (14)
$P_{G^{(j)}}(\Lambda^{(j)}(n)\leq A\;\mbox{some $n$,}\;\Lambda^{(j)}(n^{\prime})<B\;\mbox{all $n^{\prime}<n$})$ $\leq\beta$ (15)
for desired type I and II error probabilities $\alpha$ and $\beta$. The most common way of choosing the critical values $A,B$ is to use the closed-form Wald-approximations
$A=A_{W}(\alpha,\beta):=\log\left(\frac{\beta}{1-\alpha}\right)+\rho,\quad B=B_{W}(\alpha,\beta):=\log\left(\frac{1-\beta}{\alpha}\right)-\rho$ (16)
[…]
In order to apply the above to the $j$ the stream in the multiple testing set up for given step value vectors $\bm{\alpha}$ and $\bm{\beta}$, a slight extension of Bartroff and Song, (2020, Theorem 5.1) shows that choosing
$A_{k}^{(j)}=A_{W}(\widetilde{\alpha}_{k},\beta_{k}),\quad B_{k}^{(j)}=B_{W}(\alpha_{k},\widetilde{\beta}_{k})\quad\mbox{for}\quad k\in[J],$ (17)
where
$\widetilde{\alpha}_{k}:=\frac{\alpha_{1}(1-\beta_{k})}{1-\beta_{1}}\quad\mbox{and}\quad\widetilde{\beta}_{k}:=\frac{\beta_{1}(1-\alpha_{k})}{1-\alpha_{1}}\quad\mbox{for}\quad k\in[J],$ (18)
satisfies (4)-(5) up to Wald’s approximations, and that $\widetilde{\alpha}_{k}+\beta_{k}\leq 1$ and $\alpha_{k}+\widetilde{\beta}_{k}\leq 1$ for all $k\in[J]$. Although, strictly speaking, the inequalities (4)-(5) will only be guaranteed to hold approximately using these approximations, Hoel et al., (1971) show that the actual type I and II error probabilities can only exceed the nominal values by a negligibly small amount in the worst case, and the difference approaches 00 for small nominal values, which is relevant in the present multiple testing situation where we will utilize fractions of $\alpha$ and $\beta$. Alternatives to the Wald approximations in the simple vs. simple set up using a Monte Carlo or to replace the terms in (16) by the conservative values $\log\beta$ and $\log(1/\alpha)$, respectively, for which (14)-(15) always hold (see Hoel et al.,, 1971) and proceed similarly.
3 Procedures controlling type I and II error metrics
3.1 The Generic Sequential Step Down Procedure
Next we review the generic sequential step down procedure, defined in Bartroff, (2018). Versions of this procedure, implemented with certain step values $\bm{\alpha},\bm{\beta}$, will produce our proposed type I and II FDR and pFDR controlling sequential procedures. We assume that test statistics are implemented with $\bm{\alpha},\bm{\beta}$, and the critical values referred to are those satisfying (4)-(5).
We describe the procedure in terms of stages of sampling, between which reject/accept decisions are made. Let $\mathcal{J}_{i}\subseteq[J]$ ($i=1,2,\ldots$) denote the index set of the active data streams, those whose corresponding null hypothesis $H^{(j)}$ has been neither accepted nor rejected yet at the beginning of the $i$th stage of sampling, and $n_{i}$ denote the cumulative sample size of any active test statistic up to and including the $i$th stage. The total number of null hypotheses that have been rejected (resp. accepted) at the beginning of the $i$th stage is denoted by $r_{i}$ (resp. $c_{i}$). Accordingly, set $\mathcal{J}_{1}=[J]$, $n_{0}=0$, $r_{1}=c_{1}=0$. Let ||$\cdot$|| ⋅ | denote set cardinality. Then the $i$th stage of sampling ($i=1,2,\ldots$) of the Generic Sequential Step Down Procedure implemented with step values $\bm{\alpha},\bm{\beta}$ proceeds as follows.
  1. 1.
    Sample the active streams $\{X_{n}^{(j)}\}_{j\in\mathcal{J}_{i},\;n>n_{i-1}}$ until $n$ equals
    $n_{i}=\inf\left\{n>n_{i-1}:\widetilde{\Lambda}^{(j)}(n)\not\in(a_{c_{i}+1},b_{r_{i}+1})\quad\mbox{for some}\quad j\in\mathcal{J}_{i}\right\}.$ (19)
  2. 2.
    Order the active test statistics
    $\widetilde{\Lambda}^{(j(n_{i},1))}(n_{i})\leq\widetilde{\Lambda}^{(j(n_{i},2))}(n_{i})\leq\ldots\leq\widetilde{\Lambda}^{(j(n_{i},|\mathcal{J}_{i}|))}(n_{i}),$
    where $j(n_{i},\ell)$ denotes the index of the $\ell$th ordered active statistic at the end of stage $i$.
  3. 3.
    1. (a)
      If the upper boundary in (19) has been crossed, $\widetilde{\Lambda}^{(j)}(n_{i})\geq b_{r_{i}+1}$ for some $j\in\mathcal{J}_{i}$, then reject the $t_{i}\geq 1$ null hypotheses
      $H^{(j(n_{i},|\mathcal{J}_{i}|))},H^{(j(n_{i},|\mathcal{J}_{i}|-1))},\ldots,H^{(j(n_{i},|\mathcal{J}_{i}|-t_{i}+1))},$
      where
      $t_{i}=\max\left\{t\in[|\mathcal{J}_{i}|]:\widetilde{\Lambda}^{(j(n_{i},\ell))}(n_{i})\geq b_{r_{i}+|\mathcal{J}_{i}|-\ell+1}\quad\mbox{for all}\quad\ell=|\mathcal{J}_{i}|-t+1,\ldots,|\mathcal{J}_{i}|\right\},$
      and set $r_{i+1}=r_{i}+t_{i}$. Otherwise set $r_{i+1}=r_{i}$.
    2. (b)
      If the lower boundary in (19) was crossed, $\widetilde{\Lambda}^{(j)}(n_{i})\leq a_{c_{i}+1}$ for some $j\in\mathcal{J}_{i}$, then accept the $t_{i}^{\prime}\geq 1$ null hypotheses
      $H^{(j(n_{i},1))},H^{(j(n_{i},2))},\ldots,H^{(j(n_{i},t_{i}^{\prime}))},$
      where
      $t_{i}^{\prime}=\max\left\{t\in[|\mathcal{J}_{i}|]:\widetilde{\Lambda}^{(j(n_{i},\ell))}(n_{i})\leq a_{c_{i}+\ell}\quad\mbox{for all}\quad\ell=1,\ldots,t\right\},$
      and set $c_{i+1}=c_{i}+t_{i}^{\prime}$. Otherwise set $c_{i+1}=c_{i}$.
  4. 4.
    Stop if there are no remaining active hypotheses, $r_{i+1}+c_{i+1}=J$. Otherwise, let $\mathcal{J}_{i+1}$ be the indices of the remaining active hypotheses and continue on to stage $i+1$.
Thus the procedure samples all active data streams until at least one of the active null hypotheses can be accepted or rejected, indicated by the stopping rule (19). At that point, step down rejection/acceptance rules are used in steps 3a/3b to reject/accept some active null hypotheses. After updating the list of active hypotheses, the process is repeated until no active hypotheses remain.
3.2 Untruncated sequential procedures controlling FDR/FNR or pFDR/pFNR
For $m\in\{0,1,\ldots,J\}$ and step values $\bm{\alpha}$, define
$D(\bm{\alpha},m)$ $=m\left(\sum_{j=1}^{J-m+1}\frac{\alpha_{j}-\alpha_{j-1}}{j}+(J-m)\sum_{j=J-m+2}^{J}\frac{\alpha_{j}-\alpha_{j-1}}{j(j-1)}\right),$ (20)
$D(\bm{\alpha})$ $=\max_{0\leq m\leq J}D(\bm{\alpha},m).$ (21)
Our main results in this section, Theorems 3.1 and 3.2, show that the generic sequential step down procedure applied with step values $\bm{\alpha}$ have error metrics (e.g., FDR) bounded above by $D(\bm{\alpha},m_{0})$ where $m_{0}$ is the number of true null hypotheses, which is in turn bounded above by $D(\bm{\alpha})$. Since $m_{0}$ is usually unknown in practice, this latter bound is typically more useful in practice, and the former is primarily of interest for theoretical understanding of the procedure, although there are situations wherein the number of true nulls is assumed to be known (e.g., Song and Fellouris,, 2017; He and Bartroff,, 2021). It follows from $D(\bm{\alpha},m)$ being linear in $\bm{\alpha}$ that
$D(c\bm{\alpha})=cD(\bm{\alpha})\quad\mbox{for any $c>0$}.$ (22)
Hence an error metric of a procedure utilizing step values $\bm{\alpha}$ being bounded above by $D(\bm{\alpha})$ is equivalent to saying that, for a desired value $q\in(0,1)$ of the error metric, the procedure utilizing step values $\widetilde{\bm{\alpha}}=q\bm{\alpha}/D(\bm{\alpha})$ has the error metric bounded above by $q$.
The proofs of Theorems 3.1 and 3.2 are delayed until the Appendix.
Theorem 3.1.
The generic sequential step down procedure implemented with step values $\bm{\alpha},\bm{\beta}$ satisfies
$\mbox{FDR}(\theta)\leq D(\bm{\alpha},m_{0})\leq D(\bm{\alpha})\quad\mbox{and}\quad\mbox{FNR}(\theta)\leq D(\bm{\beta},m_{1})\leq D(\bm{\beta})\quad\mbox{for all}\quad\theta\in\Theta$ (23)
regardless of the dependence between data streams, where $D$ is as defined in (20)-(21) and $m_{0}$ and $m_{1}$ are the numbers of true and false null hypotheses, respectively. In particular, given $q_{1},q_{2}\in(0,1)$, the generic sequential step down procedure implemented with step values $\widetilde{\bm{\alpha}}=q_{1}\bm{\alpha}/D(\bm{\alpha})$, $\widetilde{\bm{\beta}}=q_{2}\bm{\beta}/D(\bm{\beta})$ satisfies
$\mbox{FDR}(\theta)\leq q_{1}\quad\mbox{and}\quad\mbox{FNR}(\theta)\leq q_{2}\quad\mbox{for all}\quad\theta\in\Theta$ (24)
regardless of the dependence between data streams.
For the next theorem, recall that $R$ denotes the number of null hypotheses rejected by the procedure in question.
Theorem 3.2.
Let $D$ be as defined in (20)-(21) and $m_{0}$ and $m_{1}$ denote the numbers of true and false null hypotheses, respectively. Fix arbitrary $\theta\in\Theta$.
  1. 1.
    If $\gamma_{1},\gamma_{2}>0$ are values such that the generic sequential step down procedure implemented with step values $\bm{\alpha},\bm{\beta}$ satisfies
    $P_{\theta}(R>0)\geq\gamma_{1}\quad\mbox{and}\quad P_{\theta}(R<J)\geq\gamma_{2},$ (25)
    then this procedure satisfies
    $\mbox{pFDR}(\theta)\leq\frac{D(\bm{\alpha},m_{0})}{P_{\theta}(R>0)}\leq\frac{D(\bm{\alpha})}{\gamma_{1}}\quad\mbox{and}\quad\mbox{pFNR}(\theta)\leq\frac{D(\bm{\beta},m_{1})}{P_{\theta}(R<J)}\leq\frac{D(\bm{\beta})}{\gamma_{2}}$ (26)
    regardless of the dependence between data streams.
  2. 2.
    In particular, given $q_{1},q_{2}\in(0,1)$, the generic sequential step down procedure implemented with step values $\widetilde{\bm{\alpha}}=q_{1}\gamma_{1}\bm{\alpha}/D(\bm{\alpha})$, $\widetilde{\bm{\beta}}=q_{2}\gamma_{2}\bm{\beta}/D(\bm{\beta})$ satisfies
    $\mbox{pFDR}(\theta)\leq q_{1}\quad\mbox{and}\quad\mbox{pFNR}(\theta)\leq q_{2}$ (27)
    regardless of the dependence between data streams, if $\gamma_{1},\gamma_{2}>0$ satisfy (25) for this procedure.
  3. 3.
    For $j\in[J]$ let
    $\gamma_{1j}$ $=P_{\theta^{(j)}}(\widetilde{\Lambda}^{(j)}(n)\geq b_{1}\;\mbox{some $n$,}\;\widetilde{\Lambda}^{(j)}(n^{\prime})>a_{J}\;\mbox{all $n^{\prime}<n$}),$ (28)
    $\gamma_{2j}$ $=P_{\theta^{(j)}}(\widetilde{\Lambda}^{(j)}(n)\leq a_{1}\;\mbox{some $n$,}\;\widetilde{\Lambda}^{(j)}(n^{\prime})<b_{J}\;\mbox{all $n^{\prime}<n$}).$
    If $\max_{j\in[J]}\gamma_{1j}>0$ and $\max_{j\in[J]}\gamma_{2j}>0$, then (25)-(26) hold with $\gamma_{i}=\max_{j\in[J]}\gamma_{ij}$, $i=1,2$.
4 Truncated, rejective procedures
In this section we describe versions of our procedures which only stop early to reject (rather than accept) null hypotheses and thus only explicitly control the corresponding type I multiple testing error rate (FDR or pFDR), recorded in Theorems 4.1 and 4.2. This setting naturally corresponds with having a streamwise maximum sample size (or “truncation point”) $\overline{N}$ which we assume throughout this section. For this reason we refer to them as “truncated, rejective” versions of the procedures. These procedures may be preferable in certain situations such as when (a) a null hypothesis being true represents the system being “in control” and therefore continued sampling (rather than stopping) is desirable, (b) there is a maximum sample size imposed on the data streams possible preventing simultaneous achievement of the nominal error bounds (4)-(5), or (c) the type II multiple testing error rate (e.g, FNR) $q_{2}$ is not well-motivated. In any of theses cases, one may prefer to drop the requirement that the type II multiple testing error rate be strictly controlled at an arbitrary level $q_{2}$ and use one of the rejective procedures which, roughly speaking, are similar but ignore the lower stopping boundaries $A_{k}^{(j)}$. On the other hand, if $q_{2}$ is not well motivated but the statistician prefers early stopping under the null hypotheses, we encourage the use of one of our procedures in Section 3 with both early stopping to reject and accept null hypotheses, and treat $q_{2}$ as a parameter to be chosen to give a procedure with other desirable operating characteristics, such as expected total or streamwise maximum sample size.
4.1 Set up and critical values
The set up for rejective procedures requires a few modifications. Let the data streams $X_{n}^{(j)}$, test statistics $\Lambda^{(j)}(n)$, and parameters $\theta^{(j)}$ and $\theta$ be as in Section 2. Since only the type I multiple testing error rate, FDR or pFDR, will be explicitly controlled we only require specification of null hypotheses $H^{(j)}\subseteq\Theta^{(j)}$ and not of alternative hypotheses $G^{(j)}$. We assume a streamwise maximum sample size $\overline{N}$ for each stream, but with only notational changes, what follows could be formulated by stream-specific truncation points $\{\overline{N}^{(j)}\}_{j\in[J]}$ or with sample sizes other than $1,\ldots,\overline{N}$.
Given step values $\bm{\alpha}$, we adapt our definition from Section 2.2 of the test statistics $\{\Lambda^{(j)}(n)\}_{j,n}$ being implemented with step values $\bm{\alpha}$ to this truncated, rejective setting if, for all $j,k\in[J]$, the critical values $B_{1}^{(j)},\ldots,B_{J}^{(j)}$ satisfy
$P_{\theta^{(j)}}\left(\Lambda^{(j)}(n)\geq B_{k}^{(j)}\;\mbox{for some $n\leq\overline{N}$}\right)\leq\alpha_{k}\quad\mbox{for all}\quad\theta^{(j)}\in H^{(j)},$
as well as (6) and (8) without loss of generality. We let the standardizing functions $\varphi^{(j)}$ be any increasing functions such that $b_{k}=\varphi^{(j)}(B_{k}^{(j)})$ does not depend on $j$, and $\widetilde{\Lambda}^{(j)}(n)=\varphi^{(j)}(\Lambda^{(j)}(n))$ denote the standardized statistics.
In the next section we give the truncated, rejective version of the generic step down procedure, and then in Theorems 4.1 and 4.2 state their FDR and pFDR controlling properties. The proofs are similar to the proofs of Theorems 3.1 and 3.2 and are sketched in the Appendix.
4.2 The generic rejective sequential step down procedure
With the notation of Section 3.1, the $i$th stage ($i=1,2,\ldots$) of the Generic Rejective Sequential Step Down Procedure with step values $\bm{\alpha}$ proceeds as follows.
  1. 1.
    Sample the active streams $\{X_{n}^{(j)}\}_{j\in\mathcal{J}_{i},\;n>n_{i-1}}$ until $n$ equals
    $n_{i}=\overline{N}\wedge\inf\left\{n>n_{i-1}:\widetilde{\Lambda}^{(j)}(n)\geq b_{r_{i}+1}\quad\mbox{for some}\quad j\in\mathcal{J}_{i}\right\}.$ (29)
  2. 2.
    If $n_{i}=\overline{N}$ and no test statistic has crossed the critical value in (29), accept all active null hypotheses and terminate the procedure. Otherwise, proceed to Step 3.
  3. 3.
    Order the active test statistics
    $\widetilde{\Lambda}^{(j(n_{i},1))}(n_{i})\leq\widetilde{\Lambda}^{(j(n_{i},2))}(n_{i})\leq\ldots\leq\widetilde{\Lambda}^{(j(n_{i},|\mathcal{J}_{i}|))}(n_{i})$
    and reject the $t_{i}\geq 1$ null hypotheses
    $H^{(j(n_{i},|\mathcal{J}_{i}|))},H^{(j(n_{i},|\mathcal{J}_{i}|-1))},\ldots,H^{(j(n_{i},|\mathcal{J}_{i}|-t_{i}+1))},$
    where
    $t_{i}=\max\left\{t\in[|\mathcal{J}_{i}|]:\widetilde{\Lambda}^{(j(n_{i},\ell))}(n_{i})\geq b_{r_{i}+|\mathcal{J}_{i}|-\ell+1}\quad\mbox{for all}\quad\ell=|\mathcal{J}_{i}|-t+1,\ldots,|\mathcal{J}_{i}|\right\}.$
  4. 4.
    If $r_{i}+t_{i}=J$ or $n_{i}=\overline{N}$, terminate the procedure. Otherwise, set $r_{i+1}=r_{i}+t_{i}$, let $\mathcal{J}_{i+1}$ be the indices of the remaining hypotheses, and continue on to stage $i+1$.
4.3 Truncated, rejective procedures controlling FDR or pFDR
Theorem 4.1.
The generic rejective sequential step down procedure implemented with step values $\bm{\alpha}$ satisfies
$\mbox{FDR}(\theta)\leq D(\bm{\alpha},m_{0})\leq D(\bm{\alpha})\quad\mbox{for all}\quad\theta\in\Theta$ (30)
regardless of the dependence between data streams, where $D$ is as defined in (20)-(21) and $m_{0}$ is the number of true null hypotheses. In particular, given $q_{1}\in(0,1)$, the generic rejective sequential step down procedure implemented with step values $\widetilde{\bm{\alpha}}=q_{1}\bm{\alpha}/D(\bm{\alpha})$ satisfies
$\mbox{FDR}(\theta)\leq q_{1}\quad\mbox{for all}\quad\theta\in\Theta$ (31)
regardless of the dependence between data streams.
Theorem 4.2.
Let $D$ be as defined in (20)-(21) and $m_{0}$ denote the numbers of true null hypotheses. Fix arbitrary $\theta\in\Theta$.
  1. 1.
    If $\gamma_{1}>0$ is such that the generic rejective sequential step down procedure implemented with step values $\bm{\alpha}$ satisfies
    $P_{\theta}(R>0)\geq\gamma_{1},$ (32)
    then this procedure satisfies
    $\mbox{pFDR}(\theta)\leq\frac{D(\bm{\alpha},m_{0})}{P_{\theta}(R>0)}\leq\frac{D(\bm{\alpha})}{\gamma_{1}}$ (33)
    regardless of the dependence between data streams.
  2. 2.
    In particular, given $q_{1}\in(0,1)$, the generic rejective sequential step down procedure implemented with step values $\widetilde{\bm{\alpha}}=q_{1}\gamma_{1}\bm{\alpha}/D(\bm{\alpha})$ satisfies
    $\mbox{pFDR}(\theta)\leq q_{1}$ (34)
    regardless of the dependence between data streams, if $\gamma_{1}>0$ satisfy (25) for this procedure.
  3. 3.
    For $j\in[J]$ let
    $\gamma_{1j}=P_{\theta^{(j)}}(\widetilde{\Lambda}^{(j)}(n)\geq b_{1}\;\mbox{some $n\leq\overline{N}$}).$ (35)
    If $\max_{j\in[J]}\gamma_{1j}>0$, then (32)-(33) hold with $\gamma_{1}=\max_{j\in[J]}\gamma_{1j}$.
5 Simulation studies
In this section we present the results of simulation studies of the sequential step down procedures described above in Sections 3 and 4 in order to evaluate their operating characteristics and compare them to analogous fixed-sample procedures. All computations in this section and the analysis of the UK Yellow Card pharmacovigilence database in Section 6 were performed using our Python package available from github.com/bartroff792/mult-seq-dependence. First we describe the distributional set ups for our simulation studies, and in then discuss the results in Tables 2 and 3 and conclude by comparing our proposed sequential procedures with appropriate fixed-sample competitors.
In order to examine the performance of the proposed sequential procedures on correlated Poisson and Binomial data we (i) generated correlated normally distributed data from a Gaussian copula (Trivedi and Zimmer,, 2007), (ii) applied in the inverse Gaussian c.d.f. to this data to obtain quantiles, and (iii) applied Poisson and Binomial (respectively) c.d.f.s to the quantiles to obtain correlated data of those marginal distributions. That is, for each time point $i$ we first generated a $J$-dimensional multivariate normal vector $Y_{i}\sim N_{J}(0,\Sigma)$, using the Toeplitz covariance matrix $\Sigma_{jj^{\prime}}=\rho^{|j-j^{\prime}|}$ for all $j,j^{\prime}\in[J]$ and where $\rho$ is a chosen value. Then the $J$-dimensional vector of quantiles $Q_{i}=\Phi^{-1}(Y_{i})$ was computed where $\Phi$ is the standard normal c.d.f. Finally the data values $X_{i}^{(j)}$, $j\in[J]$, were set to be
$X_{i}^{(j)}=\begin{cases}\max\left\{n:\;F_{\lambda_{j}}(n)\leq Q_{ij}\right\},&\mbox{ for the Poisson case},\\ \bm{1}\{Q_{ij}\leq p_{j}\},&\mbox{for the $\mbox{Binomial}$ case.}\end{cases}$
Here $F_{\lambda}$ is the c.d.f. of the Poisson distribution with mean $\lambda$, and the $\lambda_{j}$ and $p_{j}$ are the specified means of the data in the $j$th stream in the Poisson and Binomial cases, respectively. This process was repeated independently at each needed time point $i$. The specified value $\rho$ above equals correlation of “adjacent” elements of $Y_{i}$, corresponding to adjacent data streams, and we note that choosing $\rho<0$, as we do in some of our simulations below, leads to negative dependence among data streams and hence is outside the PRDS condition under which step up and down procedures are already well known (e.g., Finner et al.,, 2009) to control FDR.
[…]
$m_{0}$ $N_{FSS}$ $EN_{Seq}$ SE FDR SE FNR SE
0 101 36.0 0.34 0.000 0.000 0.111 0.010
1 105 39.8 0.34 0.009 0.001 0.079 0.006
3 101 45.9 0.33 0.027 0.002 0.049 0.004
5 97 50.5 0.32 0.047 0.003 0.031 0.002
7 103 53.9 0.33 0.069 0.004 0.020 0.002
9 113 55.1 0.32 0.109 0.007 0.007 0.001
10 - 55.2 0.32 0.168 0.012 0.000 0.000
[…]
$m_{0}$ $N_{FSS}$ $EN_{Seq}$ SE FDR SE FNR SE
0 83 31.6 0.22 0.000 0.000 0.107 0.010
1 79 34.3 0.23 0.009 0.001 0.077 0.006
3 77 38.1 0.24 0.029 0.002 0.052 0.004
5 73 40.4 0.25 0.050 0.003 0.038 0.003
7 79 41.6 0.26 0.077 0.004 0.023 0.002
9 99 41.0 0.26 0.119 0.008 0.007 0.001
10 - 40.1 0.23 0.172 0.012 0.000 0.000
[…]
6 Application: The UK’s Yellow Card Scheme pharmacovigilance database
6.1 Introduction
The United Kingdom’s Yellow Card Pharmacovigilence Database (yellowcard.mhra.gov.uk/information), run by the Medicines and Healthcare Products Regulatory Agency (MHRA), collects voluntary reports on safety and side-effects from the public and healthcare professionals in the UK on a host of healthcare treatments including medicines, vaccines, medical devices, and even e-cigarettes; in what follows, for simplicity we refer to the treatments in Yellow Card generically as “drugs.” The Yellow Card data collection scheme, which now includes a mobile phone app, began in 1964, spurred by the thalidomide crisis. Its use has grown steadily since then, now receiving more than 20,000 reports of possible side effects each year, and totaling more than half a million reports in the scheme’s first 40 years.
We created a Python script to download and analyze “Interactive Drug Analysis Profiles” PDF reports from Yellow Card on all of the roughly 2,800 different drugs in the database through February 2016. Figure 1 is a heatmap showing the number of drugs (indicated by color) in these reports as a function of their number of reported adverse action reactions ($y$ axis, on the $\log_{10}$ scale) and number of years ($x$ axis) the drug has been in the database, calculated from the starting date for collection of reports on the drug until the closure date of the drug’s most recent summary report. The figure shows that some of the drugs have many thousands of reaction reports, collected over decades. However, the majority of drug entries have fewer reports, collected over less than one decade, emphasizing the need for nimble, sequential monitoring of this data.
Figure 1: Number of drugs (indicated by heatmap color) in the Yellow Card data base in February 2016 as a function of total adverse reaction reports ($y$ axis, on the $\log_{10}$ scale) and the number of years ($x$ axis) the drug has been in the database.
6.2 Illustrating the sequential procedures’ performance on the database
Some Yellow Card data has been considered in the statistics literature on multiple testing (e.g., Döhler,, 2018), but the inherently streaming nature of the data has not been taken into account, with only fixed-sample multiple testing methods being applied to it so far, to our knowledge. In order to demonstrate the behavior of the sequential procedures proposed above on data like that in Yellow Card, we focus on a particular type of side effect report, amnesia, in the Yellow Card reports.
The MHRA informed us at the time of our data download that the exact time series of the reports to Yellow Card were not being made available, so it is not possible to apply our sequential procedures to the data precisely as it was being received by Yellow Card. But, for each of the drug reports downloaded as described above and illustrated in Figure 1, we were able to obtain the average number of amnesia and other side effect reports per year, the starting date for collection of reports for each drug, and the date of the closure of the drug’s most recent summary report. So here we apply our sequential procedures to data streams simulated using the actual yearly average rates obtained from Yellow Card. This approach, similar to the parametric Bootstrap (Efron and Tibshirani,, 1994), makes our simulation as close as possible to real Yellow Card data, given the limitations in data availability.
The Yellow Card data was used to simulate future drug reports as follows. At each time step a Poisson number of amnesia and non-amnesia reports for each drug was generated as correlated Poisson processes with rates
$\lambda_{amn}^{(j)}$ =(|$=\left(|\left\{\text{amnesia reports from drug~{}$j$}\right\}|+1\right)/T_{j}$ (36)
[…]
$T_{j}$ $=(\mbox{drug~{}$j$ most recent report})-(\mbox{drug~{}$j$ start date}).$
Note that we have added $1$ amnesia and non-amnesia report to each drug in (36)-(37) to account for issues with rare or exotic drugs like inherently higher variance in use and reporting. These rates are visualized in Figure 2, where the blue, green, and red points are the drugs with “low,” “medium,” and “high” amnesia report rates according to (38) and(39).
Figure 2: Drug side effect rates (36)-(37) from the Yellow Card database data. Blue points are those drugs $j$ with $p^{(j)}\leq p_{H}$, green those with $p_{H}<p^{(j)}<p_{G}$, and red those with $p^{(j)}\geq p_{G}$, as defined by (38) and (39).
In order to generate a medically realistic correlation structure for the drugs, we utilized BioGPT (Luo et al.,, 2022), a generative language model trained on biomedical research articles. For each drug under consideration, we calculated the element-wise mean of the BioGPT’s token embedding for “[DrugName] as a drug” and then applied $K$-means clustering to obtain $K=15$ clusters of the drugs. Within each cluster, the copula method described in Section 5 was then used to generate the Poisson data using the Toeplitz covariance matrix given there, but with the correlation parameter $\rho$ generated as a Beta$(4,2)$ random variable, shifted and scaled to have support $[-1,1]$. In this way, drugs that were medically related according to BioGPT were more likely to receive non-trivial correlation.
In order to account accurately for the differing number of reports that drugs with different prescription and use rates in the database have, we set up the hypotheses tests not in terms of the Poisson rate parameters (36)-(37) but rather a scaled parameter
$p^{(j)}=\frac{\lambda_{amn}^{(j)}}{\lambda_{amn}^{(j)}+\lambda_{-amn}^{(j)}}\approx\frac{\mbox{\# amnesia reports for drug~{}$j$}}{\mbox{total \# reports for drug~{}$j$}}.$ (38)
The data generation scheme above is equivalent to a Binomial number of amnesia reports being generated at each time step, with success probability (38) and number of trials equal to the random total number of Poisson reports.
To set up our hypothesis tests, we utilize the database to calculate
$p_{H}=\mbox{50th percentile of}\;\{p^{(j)}\},\quad\mbox{and}\quad p_{G}=\mbox{90th percentile of}\;\{p^{(j)}\}$ (39)
which we use as surrogates for “typical” and “extreme” Binomial rates of amnesia reports, respectively. After calculating $p_{H}$ and $p_{G}$ on the entire Yellow Card database, we then applied out sequential procedure to the $J=1800$ drugs with the highest total side effect reports in the database. This reduction still maintains a large number of data streams but also allows us to filter out drugs that have only recently been made available with too small a number of reports, but also the opposite extreme of filtering out obscure and antiquated drugs that exist in the database (e.g., “Mammalian Blood” and “Wool Fat”).
For each of these $J=1800$ drugs we tested the hypotheses
$H^{(j)}:p^{(j)}\leq p_{H}\quad\mbox{vs.}\quad G^{(j)}:p^{(j)}\geq p_{G}>p_{H}$
representing that the drug has a “typical” versus abnormally high rate of amnesia reports. For these we implemented the open-ended sequential test described in Theorem 3.1 for Binomial data using FDR controlled at $q_{1}=0.05$ and FNR at $q_{2}=0.15$ using with the Wald approximations (17)-(18) the BH step values (12). Running this data generation scheme with $1800$ data streams a single time is computationally intensive even on a Google Cloud Computing instance, so rather than report Monte Carlo statistics, in Table 4 we give the output of a typical run and there list the first 15 drugs whose streams terminated accept, and the first 15 to reject, the null hypothesis $H^{(j)}$. Since the simulation is set up where each “observation” is the number of reports in a year, in the table the Termination Step corresponds to the number of years required for termination of each stream. The Termination Level is a measure of the strength of the terminal action, smaller meaning a stronger acceptance/rejection of the null, and equals the cumulative value of the level $\ell$ in step 3a or 3b of the procedure as defined in Section 3.1.
While we draw no medical conclusions from our simulations, we note that some of the drugs the sequential procedure stopped to reject their null hypotheses in favor of an abnormally high level of amnesia reports, have been associated with amnesia in the biomedical literature (e.g., Chavant et al.,, 2011).
[…]
Drug Terminal Action Termination Step Termination Level
Bupropion Accept $H^{(j)}$ 1 2
Clozapine Accept $H^{(j)}$ 1 2
Etanercept Accept $H^{(j)}$ 2 8
Venlafaxine Accept $H^{(j)}$ 2 8
Varenicline Accept $H^{(j)}$ 2 8
Paroxetine Accept $H^{(j)}$ 2 8
Clostridium Tetani Accept $H^{(j)}$ 2 8
Telaprevir Accept $H^{(j)}$ 2 8
Adalimumab Accept $H^{(j)}$ 3 10
Rofecoxib Accept $H^{(j)}$ 3 10
Ranibizumab Accept $H^{(j)}$ 4 21
Amoxycillin Accept $H^{(j)}$ 4 21
Oseltamivir Accept $H^{(j)}$ 4 21
Trimethoprim Accept $H^{(j)}$ 4 21
Ciprofloxacin Accept $H^{(j)}$ 4 21
Zopiclone Reject $H^{(j)}$ 7 1
Simvastatin Reject $H^{(j)}$ 9 2
Zolpidem Reject $H^{(j)}$ 13 3
Bifonazole Reject $H^{(j)}$ 15 5
Rimonabant Reject $H^{(j)}$ 15 5
Enzalutamide Reject $H^{(j)}$ 25 7
Gabapentin Reject $H^{(j)}$ 25 7
Atorvastatin Reject $H^{(j)}$ 31 8
cAC10-vcMMAE Reject $H^{(j)}$ 32 9
Dutasteride Reject $H^{(j)}$ 33 10
Mitotane Reject $H^{(j)}$ 36 11
Boceprevir Reject $H^{(j)}$ 39 12
Lurasidone Reject $H^{(j)}$ 42 14
Asenapine Reject $H^{(j)}$ 42 14
Retigabine Reject $H^{(j)}$ 44 16
7 Conclusion
We have proposed both open-ended and truncated sequential multiple testing procedures which can control FDR or pFDR (and their type II analogs, FNR or pFNR in the open-ended case) under arbitrary dependence between the data streams. These procedures have shown large savings in average sample size compared to their fixed-sample equivalents in our simulation studies, and in Section 6 we demonstrate how these procedures may be used to monitor data streams like those coming from a pharmacovigilance database, like the UK’s Yellow Card database.
Compliance with Ethical Standards
The authors have no conflicts of interest to report.
Appendix
Appendix A Proof of Theorem 3.1
We first prove the inequalities for FDR in (23) for the sequential step down procedure with step values $\bm{\alpha},\bm{\beta}$, and then discuss those for FNR. The inequalities in (24) then follow, using (22). Our proof largely follows the arguments of Guo and Rao, (2008), with a few important differences to account for the sequential nature of our procedures.
Fix arbitrary $\theta$ and omit it from the notation, and below let $P(\cdot)$ and $E(\cdot)$ denote the probability and expectation under $\theta$ and an arbitrary joint distribution between the data streams. Without loss of generality assume $H^{(1)},\ldots,H^{(m_{0})}$ are the true hypotheses for some $m_{0}\in[J]$, the $m_{0}=0$ case being trivial since $\mbox{FDR}=0$ in this case. For $i\in[J]$ let $\tau_{i}$ denote the time at which the $i$th stream $\widetilde{\Lambda}^{(i)}$ terminates, and for $i,j\in[J]$ let
$F_{ij}=\{\widetilde{\Lambda}^{(i)}(\tau_{i})\in[b_{j},b_{j-1})\},$ (40)
setting $b_{0}=\infty$ to handle the $j=1$ case. The event $F_{ij}$ is similar (in fact, contained in) the event in (10), but it specifies which interval $[b_{j},b_{j-1})$ the $i$th statistic $\widetilde{\Lambda}^{(i)}$ is in when stopping to reject its corresponding null $H^{(i)}$; we will refer to this as $H^{(i)}$ being rejected at the $j$th level. Recalling that $R$ and $V$ denote the number of null and true null hypotheses, respectively, rejected by the procedure, define
$p_{ijk}=P(F_{ij}\cap\{R=k\})\quad\mbox{for}\quad i,j,k\in[J],$ (41)
with which we can write
$\mbox{FDR}=\sum_{i=1}^{m_{0}}\sum_{j=1}^{J}\sum_{k=j}^{J}\frac{p_{ijk}}{k},$ (42)
which is analogous to similar expressions obtained for FDR in the fixed-sample setting obtained by Benjamini and Yekutieli, (2001) and Sarkar, (2002). To see why (42) holds here for the sequential procedure, write
$\mbox{FDR}=E\left(\frac{V}{R\vee 1}\right)=\sum_{k=1}^{J}\frac{1}{k}E(V\bm{1}\{R=k\})=\sum_{k=1}^{J}\sum_{i=1}^{m_{0}}\frac{1}{k}P(\{\mbox{$H^{(i)}$ rejected}\}\cap\{R=k\})\\ =\sum_{k=1}^{J}\sum_{i=1}^{m_{0}}\sum_{j=1}^{k}\frac{1}{k}P(F_{ij}\cap\{R=k\})=\sum_{i=1}^{m_{0}}\sum_{j=1}^{J}\sum_{k=j}^{J}\frac{1}{k}p_{ijk},$
where in the last equality we have reordered the sums and used the definition of $p_{ijk}$.
We further extend the notation to expand (42). Given $k\in[J]$, $\ell\in[k\wedge m_{0}]$, and a pair of vectors $(\bm{i},\bm{j})$ in
$\Omega_{\ell k}=\{(\bm{i},\bm{j})\in[m_{0}]^{\ell}\times[k]^{\ell}:1\leq i_{i}<i_{2}<\ldots<i_{\ell}\leq m_{0}\},$
define
$p_{\bm{i}\bm{j}k}=P\left(\bigcap_{d=1}^{\ell}F_{i_{d}j_{d}}\cap\{R=k\}\cap\{V=\ell\}\right),$ (43)
which is the probability of $k$ rejections, of which the false rejections are $H^{(i_{d})}$ being rejected at the $j_{d}$th level, $d\in[\ell]$. Let
$\Omega_{\ell k}(i,j)=\{(\bm{i},\bm{j})\in\Omega_{\ell k}:(i,j)=(i_{d},j_{d})\;\mbox{for some $d\in[\ell]$}\}$
be the set of those vector pairs that include $H^{(i)}$ being rejected at the $j$th level. With these definitions we can further expand $p_{ijk}$ in the form
$p_{ijk}=\sum_{\ell=1}^{k\wedge m_{0}}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}p_{\bm{i}\bm{j}k}.$ (44)
This equality holds because, for distinct pairs $(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)$, the events in (43) correspond to different nulls being rejected, or at different levels, and are hence disjoint, thus
$\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}p_{\bm{i}\bm{j}k}=P\left(F_{ij}\cap\{R=k\}\cap\{V=\ell\}\right).$
For distinct values of $\ell$, the events in this last are clearly disjoint, hence summing over $\ell$ yields
$\sum_{\ell=1}^{k\wedge m_{0}}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}p_{\bm{i}\bm{j}k}=\sum_{\ell=1}^{k\wedge m_{0}}P\left(F_{ij}\cap\{R=k\}\cap\{V=\ell\}\right)=P(F_{ij}\cap\{R=k\})=p_{ijk}.$
With (44) established we can substitute it into (42) to obtain
$\mbox{FDR}=\sum_{i=1}^{m_{0}}\sum_{j=1}^{J}\sum_{k=j}^{J}\sum_{\ell=1}^{k\wedge m_{0}}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}\frac{p_{\bm{i}\bm{j}k}}{k}=\sum_{\ell=1}^{m_{0}}\sum_{k=\ell}^{J}\sum_{i=1}^{m_{0}}\sum_{j=1}^{k}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}\frac{p_{\bm{i}\bm{j}k}}{k}\\ =\sum_{\ell=1}^{m_{0}}\sum_{k=\ell}^{J}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}}\frac{\ell p_{\bm{i}\bm{j}k}}{k}.$ (45)
The second equality in (45) is obtained by reordering the summations, and the last equality uses
$\sum_{i=1}^{m_{0}}\sum_{j=1}^{k}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}\frac{p_{\bm{i}\bm{j}k}}{k}=\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}}\frac{\ell p_{\bm{i}\bm{j}k}}{k},$
since each pair $(\bm{i},\bm{j})\in\Omega_{\ell k}$ appears in the $\ell$ sets $\Omega_{\ell k}(i_{1},j_{1}),\ldots,\Omega_{\ell k}(i_{\ell},j_{\ell})$, and only those sets.
With the expression (45) for FDR established, we now consider constraints that the $p_{\bm{i}\bm{j}k}$ must satisfy. For the first constraint, take arbitrary $k\in[J]$, $\ell\in[k\wedge m_{0}]$, $(\bm{i},\bm{j})\in\Omega_{\ell k}$, and let $j_{(1)}\leq\ldots\leq j_{(\ell)}$ denote an ordering of the values in $\bm{j}=(j_{1},\ldots,j_{\ell})$. For an arbitrary fixed-sample step down procedure with $p_{\bm{i}\bm{j}k}$ defined analogously, Guo and Rao, (2008, Lemma 3.3) show that
$p_{\bm{i}\bm{j}k}=0\quad\mbox{if}\quad k-\ell>m-m_{0}\quad\mbox{or}\quad j_{(d)}>k-\ell+d\;\mbox{for some $d\in[\ell]$.}$ (46)
Their proof depends only on the step down structure and the values that $R,V$ can take with positive probability, and thus hold for our sequential step down procedure as well, so we do not repeat the proof here.
The other constraint relates the $p_{\bm{i}\bm{j}k}$ to the error probabilities in (10). First observe that, for any $i\in[m_{0}]$, the events $F_{i1},F_{i2},\ldots$ are disjoint, as are $\{R=1\},\{R=2\},\ldots$. Then, for any $s\in[J]$, we write (10) as
$\alpha_{s}\geq P\left(\bigcup_{j=1}^{s}F_{ij}\right)=\sum_{j=1}^{s}P(F_{ij})=\sum_{j=1}^{s}P\left(F_{ij}\cap\bigcup_{k=1}^{J}\{R=k\}\right)=\sum_{j=1}^{s}\sum_{k=1}^{J}P\left(F_{ij}\cap\{R=k\}\right)\\ =\sum_{j=1}^{s}\sum_{k=1}^{J}p_{ijk}=\sum_{j=1}^{s}\sum_{k=1}^{J}\sum_{\ell=1}^{k\wedge m_{0}}\sum_{(\bm{i},\bm{j})\in\Omega_{\ell k}(i,j)}p_{\bm{i}\bm{j}k},$ (47)
using (44) for this last equality.
Combining (45), (46), and (47), the goal of finding the worst-case joint distribution that maximizes FDR can be stated as finding the $\{p_{\bm{i}\bm{j}k}\}$ that
[…]
At this point the $\{p_{\bm{i}\bm{j}k}\}$ can be completely divorced from their original meaning about a multiple testing procedure and treated as arbitrary variables in the constrained optimization problem (48). Guo and Rao, (2008) solve a similar problem, identical to (48) except for the second constraint; their second constraint is sufficient but not necessary for our second constraint (47); about this, see Remark A.1. Guo and Rao proceed to solve this problem by producing an upper bound on FDR that coincides with FDR at its maximum $D(\bm{\alpha},m_{0})$, and then show that the maximizer satisfies both constraints. Since their second constraint is sufficient for ours, their proof applies here, and we do not repeat more of the details here. This establishes the inequalities for FDR in (23).
The inequalities for FNR in (23) are established in a completely similar way after reversing the roles of the “type I” and “type II” objects, e.g., substituting FNR for FDR, $\bm{\beta}$ for $\bm{\alpha}$, the bound (11) for (10), etc. The details are straightforward and therefore omitted here. ∎
Remark A.1.
Guo and Rao’s (2008, expression (19)) second constraint involves bounding from above the probability that a $p$-value for a true null falls in the interval $[\alpha_{j-1},\alpha_{j})$ by $\alpha_{j}-\alpha_{j-1}$. Validity of the $p$-value is not sufficient for this to hold and, for example, it may fail for valid but discrete $p$-values. Therefore their proof actually requires a condition stronger than validity on the $p$-values, for example having the exact uniform $(0,1)$ null distribution. We avoid the need for such a stronger condition by summing over $j$ in (47), and thus only need (10) which is the analog of validity for $p$-values in the fixed sample set up.
Appendix B Sketch of proof of Theorem 3.2
The proof of the inequalities for pFDR in (26) is similar to the proof of Theorem 3.1 after replacing $p_{ijk}$ in (41) by
$p_{ijk}=P(F_{ij}\cap\{R=k\}|R>0)$
and using (25) to bound $P(R>0)$ from below, leading to the factor of $1/\gamma_{1}$ in the bound. The proof of the inequalities for pFNR in (26) proceeds similarly, conditioning on $R<J$ and using $\gamma_{2}$. Part 2 of Theorem 3.2 then follows using (22). In Part 3, the event in (28) implies rejection of $H^{(j)}$, hence $\max_{j}\gamma_{1j}\leq P(R>0)$, with analogous statements applying to the type II version.
Appendix C Sketch of proofs of Theorems 4.1 and 4.2
The proof of Theorem 4.1 follows the proof of FDR control in Theorem 3.1 with $p_{ijk}$ defined in (41) but with
$F_{ij}=\{\widetilde{\Lambda}^{(i)}(\tau_{i})\in[b_{j},b_{j-1}),\quad\tau_{i}<\overline{N}\}$ (49)
rather than (40).
For Theorem 4.2, the proof proceeds the similarly but with $p_{ijk}$ replaced by
$p_{ijk}=P(F_{ij}\cap\{R=k\}|R>0),$
again with $F_{ij}$ given by (49). Then (32) is used to bound $P(R>0)$ from below, giving the factor of $1/\gamma_{1}$ in the bound. Part 2 follows from (22), and the event in (35) of Part 3 implies rejection of $H^{(j)}$, hence $\max_{j}\gamma_{1j}\leq P(R>0)$.